# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ 6.1 Rule-based Heuristic for SLID Detection
6.2 Heuristic Evaluation (Small Scale)
We apply different datasets to evaluate our heuristic.
Accuracy on ground truth datasets. Before applying our heuristic to Ethereum liquidity pools (provided in §7), we first utilize our initial ground truth to evaluate the accuracy of the proposed rule-based heuristic. The rule-based heuristic has successfully flagged the 71 liquidity pools as SLID pools and correctly labeled the others as non-SLID scams. Evaluation results demonstrate that our heuristic achieved perfect accuracy (100%, experimental figures thus ignored) when tested on the RQ1 ground truth, resulting in no false positives. This lack of false positives in this evaluation is due to the limited ground truth available in this study. Since the heuristic is tailored directly from the ground truth, there is no other ground truth for false positive determination.
Heuristic performance on legitimate liquidity pools. To ensure that the heuristic generalizes well beyond our collected ground truth dataset and remains effective for identifying liquidity pools outside this dataset, we further validate its accuracy by testing it on external pools.
Given the absence of prior research or an established ground truth for SLID liquidity pool data, we evaluate the heuristic’s ’global accuracy’ by manually analyzing liquidity pools outside of our initial dataset. Specifically, we selected 100 randomly sampled legitimate liquidity pools deployed after 2022, consistent with the time threshold used in collecting rug pull reports (§5.1). We analyze the heuristic’s true negative and false positive rates (ability to correctly identify legitimate pools as non SLID without flagging them) to provide a robust measure of its performance.
With the heuristic, none of the 100 legitimate LPs were flagged as SLID scams. They all triggered the condition of burning initial LP tokens (defined in Owner Activity Validator). However, if we remove the LP burning condition from the heuristic, eight pools were flagged as SLID. These pools exhibited owner actions of gradually taking profits, similar to those seen in SLID scams. Notably, these actions were announced in advance through tokenomics, indicating a pre-declared intention rather than fraudulent behavior. This indicates that our heuristic still keeps 100% accuracy.
@@ 6.2 Heuristic Evaluation (Small Scale)
Findings of RQ2: The rule-based heuristic for SLID liquidity pool detection, with the validator for honeypot, profit and owner-action, demonstrates to be an effective detector. When tested against the initial ground truth and a dataset of 100 legitimate liquidity pools, it achieves 100% accuracy.
@@
7 Detecting Ethereum’s SLID Scams (RQ3)
@@
We conduct a large-scale scam detection analysis (Figure 5). We apply our rule-based heuristic (§6.1) on the datasets (including liquidity pools, and user activities) collected from six leading Ethereum DEXs (§7.17.2). In addition, we propose an ML-based approach (§7.3), designed to provide earlier detection of SLID scams. This is crucial for protecting investors from scams that unfold gradually over time.
@@
Figure 5: Overall processes of our SLID detection.
7.1 Data Collection
@@
We collected longitudinal Ethereum’s LP data spanning over 67 months. We preprocessed the data to extract relevant features. We utilized both heuristic-based and machine learning algorithms to classify the pools as SLID or non-SLID .
@@
Liquidity pool information.  We first collected all the existing liquidity pools in Ethereum (denoted as PoolData) to be used for both detection methods. We utilized the well-known Geth client node [49] to retrieve all Ethereum block information, starting from block 6,627,000 (November 2, 2018, the date when the first liquidity pool was deployed) to block 20,081,867 (June 13, 2024, the time of writing this work). Then, we filtered only pools containing at least one well-established token (such as ETH, USDT, or USDC) as the pool’s legitimate base token. We avoid adding less-recognized and less-active pairs to the dataset.
@@
As a result, we have gathered a total of 319,166 deployed liquidity pools across the six largest DEXs on Ethereum [50] (Table 1), inclduing Uniswap [33], Curve Finance [12], Balancer [51], SushiSwap [13], PancakeSwap [52], and BancorSwap [53]. We extracted key features (Appendix G) for each harvested liquidity pool, including creation timestamps, token addresses involved in the pool, owner addresses, and the initial pool value in USD, using Chainalysis API [54].
Table 1: Collected data on Ethereum.

### v3

@@ III-B More LPs Exhibiting SLID Patterns
IV Common Features of SLID Scams (RQ1)
Based on examples, we developed a multi-layered filtering mechanism (§IV-A) to initiate our pilot study and find the scams with similar features (i.e., SLID-like LPs, §IV-B).
IV-A Identification of SLID LPs
We began by collecting publicly available DeFi scam reports from a broad range of sources dating back to 2022. These sources of reports included well-known platforms such as Chainabuse [41], De.fi [42], Chainalysis [43], and prominent scam warning accounts on X (formerly Twitter), including ZachXBT [44] and Pocket Universe [45]. From this collection, we identified 921 reported LPs that formed our initial dataset (can be accessed in [46]).
We then developed and applied a four-layer filtering system to isolate LPs that exhibited SLID-like characteristics:
  • $\bullet$
    Owner profit-checking layer. The layer computes the realized profit of the LP owner based on their profit-taking action from trading history and then excludes LPs in which the owner has a negative realized profit (Appendix A-C). The layer ensures accuracy by filtering out the falsified reports where tokens were not fraudulent but reported as a scam. As a result, 898 LPs bring profits to the owner out of 921 reported LPs, and 23 LPs were excluded.
  • $\bullet$
    Honeypot layer. The detection against honeypot is straightforward in LP’s deployment phase. We filtered out instances that contain honeypot vulnerability by utilising GoPlus [47] and De.Fi [42] services. We excluded 139 honeypot LPs out of 898 LPs. The remaining 759 LPs were eligible.
  • $\bullet$
    Mislabelled LPs filtering layer. This layer processes the LP instances and cross-references them with the rug pull detection methods (i.e., [24, 39]), to determine if any of the collected scams were rug pull and filters them out from the SLID-like LPs. Out of the 759 input instances, 665 instances were recognized as rug pull and filtered out, while 94 LPs moved to the next layer.
  • $\bullet$
    Owner action layer. Amongst the 94 remaining instances, we recognized that they may contain various other undetermined DeFi frauds that are irrelevant to SLID. We thus selected LPs (i) without the initial LP burned and (ii) having at least three DEX activities conducted by owners, including deposit, withdraw, buy, or sell (aligning with our observations in §1). As a result, we excluded 23 LPs and identified 71 LPs that passed all four layers.
@@ IV-A Identification of SLID LPs
IV-B Finding and Refining Features from SLID-like LPs by Comparing with Rug Pulls
To identify distinguishing features, we conducted an in-depth analysis and benchmarked the 71 filtered LPs against rug pull tokens that had been excluded mislabeled LPs filtering layerIV-A). These rug pulls were reintroduced as a behavioral baseline for comparison, as they exhibit two known fraudulent profit-taking strategies: (i) inflated selling and (ii) initial liquidity pool (LP) withdrawal.
Our analyses identify three defining properties (P1, P2, P3) commonly observed in SLID LPs.
Figure 3: 30-day liveliness of rug pull and SLID LPs.
IV-B1 Post-month liveliness analysis
Liveness is one key metric to detect scam tokens [39][48][49]. Nearly 60% of the rug pull LPs ended up “inactive” after 1 day of deployment [39].
We observed LP lifetimes for both scam types over an extended period of one month, called post-month liveliness. We also compared the average changes in user activity for LPs that remain “active” after the first month to distinguish SLID scams from rug pulls.
Our results revealed that 72.85% of SLID-like LPs remain active after the first month of deployment, whereas only 10.8% of rug pull LPs persist beyond this threshold. On the first day of the second month, trading activities for the remaining active SLID LPs were stable, showing an average reduction of 38.16% compared to the LP’s previous peak day with the highest recorded activity count. In contrast, the remaining rug pull LPs experienced a sharp decrease (95.82%) in recorded activities on the first day of the second month.
The findings led to our conclusion:
  • P1.
    The SLID LPs are more adept at remaining undetected due to their prolonged liveliness. They remain active over a longer duration compared to rug pull despite using similar LP initiation and profit-taking mechanisms, making it nearly impossible to detect based solely on liveliness status.
IV-B2 Scammer activity analysis
For both rug pulls and SLID scams, scammers were LPs owners and relied on two strategies: (i) withdrawing liquidity from the LP and (ii) selling the inflated tokens they held externally to drain the LP. We then analyzed the scammer’s activity patterns
Our results show that rug pull scammers prioritize immediate profit-taking, with 99.7% (663/665) of LPs being entirely drained within the first day, and 100% drained within the first three days. The LP owners, who are also scammers, perform 75.38 DEX activities to their LP and 15.79 profit-taking actions on average during the first month. Among these profit-taking actions, each LP includes a single, large-amount profit-taking transaction (either selling or withdrawing) that constitutes over 95% of the total LP value. On average, those scammers extract 99.33% of the LP’s value after the first month, leaving only 0.67% as unrealized profit.
In contrast to rug pulls, scammers in SLID LPs operate more gradually, conducting multiple profit-taking activities with smaller proportions. None of the SLID LPs were completely drained within the first three days after deployment, unlike 100% of rug pull LPs, which were fully drained in that same timeframe. On average, SLID scammers executed 996.59 total orders during the LP’s first month, over 10 times more than rug pull LP owners. They also carried out a mean of 423 profit-taking actions, approximately 26 times more frequent than rug pull scammers. Additionally, SLID scammers keep their profit-taking within a narrower range, with drained values varying from 7.39% to 42.93% of the LP’s total value. This controlled approach can evade detection by standard rug pull detectors.
To better demonstrate differences, we further provided two observed example tokens (Figure 4): MIND [50] (rug pull) and EMILY [51] (SLID-like). While a single profit-draining action by the MIND owner completely emptied LP and reduced the token’s price to zero, the EMILY owner conducted multiple small-proportion exits that do not heavily affect the token’s price. Although both scam types initiate fraud similarly by creating LPs and profiting through draining them via sell or withdraw orders, we observed that variations in execution duration and order proportions lead to distinct strategies for each scam type. We thus conclude that
  • P2.
    Although SLID scams employ a mechanism similar to that of rug pull, they pursue an opposite strategy characterised by a gradual approach, avoiding recognition. Scammers prioritise multiple small-proportion and slow-paced profit extraction activities, making it difficult to detect, either through observation or by using detectors.
Figure 4: Token prices of EMILY (SLID-like) and MIND (rug pull) were affected by scammer’s profit-taking activities in the first 3 days.
IV-B3 Profit analysis
The SLID strategy is intuitively expected to yield a better profit compared to rug pull by allowing scammers to conceal activities while gradually attracting more victims. We validate this speculation via a profit-changing analysis between the two scam types. Technically, we focus on both realized and unrealized profits of the LP owner.
Our results showed that the realized profits (explained in §II). SLID scams demonstrate a superior profit margin: in the first month, while rug pull scammers achieve an average realized profit of 7.5 times their investment through a single, short-term draining action, SLID LPs attain a massively higher profit margin of 10.3 times their investment through multiple small draining activities. As for unrealized profits, SLID scams also perform a notable gap in profit advantage. Rug pull scammers have nearly zero future profit while completely emptying their LPs, retaining only an average of 0.67% of LP value. In contrast, the SLID LP owner’s drainable assets in the LP after the first month retain, on average, 156% worth of their initial investment. This allows the scammer to either fully drain the unrealized profit, resulting in an additional 1.56 times profit on top of the realized profit (already taken), or to refrain from withdrawing and let LP continue to grow.
We found that, by performing a vast number of small-proportion profit-taking orders (§IV-B2), SLID scammers manage to attract more victims, resulting in a larger netflow of funds poured into LPs. Although they only take part of their possible profit in the first month, these scammers already achieve better profit margins than rug pull scammers. Furthermore, a significant portion of assets that remain in the SLID LPs retain substantial profit-scaling potential as the LP continues to grow, entrapping more future buyers.
Therefore, we conclude that,
  • P3
    SLID scams offer a higher profit margin and sustainability, providing both significant mid-term profits and future profit-scaling potential if undetected. Compared to rug pull, SLID scammers can achieve a substantial growth rate in LP value from their initial investment, making it a more lucrative and enduring scam approach.
[…]
V Rule-based SLID Detection (RQ2)
Based on our empirical insights (§IV), we step further by formally defining SLID fraudulent (§V-A) actions and introducing our heuristic (§V-B).
V-A SLID with A Formal Understanding
System model. We consider a DEX where an LP $\mathcal{P}$ (liquidity pool) allows trading between:
  • A base token, denoted $B$ (e.g., ETH, USDC), and
  • A paired token, denoted $P$ (typically a new or obscure token).
There are two main participants:
  • Owner ($O$): deploys the LP and adds initial liquidity.
  • Investors: interact with the LP through swaps, deposits, and withdrawals.
We define the following terms:
  • $t_{0}$: time of LP deployment.
  • $B_{0},P_{0}$: initial deposited amounts of base and paired tokens.
  • $p^{t}$: price of token $P$ in terms of $B$ at time $t$.
  • $B^{t}$: base token balance in the pool at time $t$.
  • $V^{t}$: trading volume at time $t$.
  • $X^{t}$, $Y^{t}$: total amount of $P$ bought/sold by users up to time $t$.
  • $W^{t}$: cumulative amount of $B$ withdrawn by $O$ by time $t$.
We assume the owner:
  • Retains full control over the initial LP tokens (denoted $\texttt{LPT}_{\text{init}}$).
  • Possesses an amount of inflated paired tokens $P$.
We identify three core fraudulent actions (rooted in §III).
Retained LP Token Control. The owner does not burn initial LP tokens: $\texttt{burn}(\texttt{LPT}_{\text{init}})=\text{false}$ This permits future liquidity withdrawals without investor approval. Inflated token selling. The owner sells their private inflated supply of $P$ over time via small-volume trades: $\mathcal{S}=\{(p^{i},q^{i})\}_{i=1}^{n},$ where each $q^{i}$ is a single trade of an owner’s $P$ token at price $p^{i}$ (at time $t_{i}$) satisfying: $q^{i}$ $<\delta\cdot B^{t_{i}},\quad(\text{small relative to pool size})$ $|p^{i+1}-p^{i}|$ $\approx 0\quad(\text{stable price})$ with $\delta\ll 1$. Incremental liquidity withdrawals. The owner performs multiple small withdrawals of base tokens: $\mathcal{W}=\{w^{j}\}_{j=1}^{m}$ such that: $\sum_{j=1}^{m}w^{j}$ $>\beta\quad(\text{significant total})$ $w^{j}$ $<\epsilon\quad(\text{each withdrawal is small})$ with $\beta\in(0,1]$, $\epsilon\ll 1$.
Formal Definition of SLID Scam.
Definition 1 (SLID)
A liquidity pool is defined as a SLID scam over the time interval $t_{0},t_{k}]$ if the following hold:
  1. 1.
    Control Retention: $\texttt{burn}(\texttt{LPT}_{\text{init}})=\text{false}$.
  2. 2.
    Inflated Sales: A sale sequence $\mathcal{S}$ exists such that:
    $\quad q^{i}<\delta\cdot B^{t_{i}}$
  3. 3.
    Withdrawals: A withdrawal sequence $\mathcal{W}$ exists such that:
    $\sum w^{j}>\beta,\quad w^{j}<\epsilon$
  4. 4.
    Stable Appearance: Price and volume exhibit low volatility:
    $\texttt{std}(p(t))<\theta_{p},\quad\texttt{std}(V(t))<\theta_{v}$
    for some small thresholds $\theta_{p}$, $\theta_{v}$.
Using the actions above, we label a pool SLID when the owner keeps control of the LP (1) and the pool still looks normal to users (4), and there is clear value extraction either through owner‑side sells (2) or owner withdrawals (3). Note that it is possible that the scammer leverages (2) and (3) in combination in an LP to exploit the pool and take profit in a SLID case.
Threat model. A SLID adversary (i.e., LP owner) aims to:
  • Stealthily extract profit through long-term operations,
  • Maintain an illusion of a healthy market (low volatility),
  • Evade on-chain scam detectors and user suspicion.
Remark. To ensure our SLID definition aligns with actual filtering outcomes, we conducted a second-round manual validation. From 71 SLID-like LPs, we randomly sampled 14 (20%) and verified the presence of all three fraudulent behaviors: unburned LP tokens (action-①), inflated token selling (action-②), and gradual withdrawal (action-③). All sampled LPs exhibited the full behavioral pattern, confirming the accuracy of our definition. Rather than enforcing rigid thresholds, our definition embraces the fuzzy, incomplete nature of real-world scams. Parameters such as lifespan (P1), profit-taking (P2), and total profit (P3) are assessed through partial or fuzzy matching, capturing variants of canonical SLID traits (see more thoughts in Appendices A-F and A-E). This flexible design reflects the strategic behavior of rational attackers, who aim to maximize profit while delaying detection, mirroring signaling game dynamics applied in deception [52, 53, 54, 55]. By emphasizing retained control, incremental drains, and price stability, our heuristic definition captures the trade-off between stealth and sustained exploitation.
V-B From Definition to Rule-based Heuristic
Following the above, we recognize that real-world scams can exhibit a wide range of behaviors (e.g., low to high profits, varying transaction frequencies, different lifetimes). To capture this variability, we designed a dynamic heuristic detection framework based on the definition and with flexible thresholds.
Our heuristic has three validators: Honeypot Validator, Profit Validator, and Owner Activity Validator. Each targets different properties from the extracted data of LPs (cf. RQ1 findings). If one LP is flagged by all three validators (i.e., exhibiting all observed traits), it will be labelled as SLID scam.
Honeypot validator.  This component assesses whether tokens are honeypots. A successful SLID scam depends on a gradual increase in victim participation and buying volume (P2).
Our honeypot validator employs rule-based criteria derived from prior research of [25, 56]. Instead of simply replicating exact detection methods, we apply criteria to specific features extracted from token’s smart contracts (implemented in [57]) and set an upper threshold to identify honeypot signals.
In particular, we extracted the following features from the paired token’s smart contract: “buy tax”, “sell tax”, “cannot buy”, “cannot sell all”, “slippage modifiable”, “transfer pausable”, “personal slippage modifiable”, and “trading cooldown”. These features were evaluated based on specific criteria; each indicates different scenarios in which the LP’s owner can illegally interfere with the user’s activity in the LP for gaining profit. Here, we provide an intuitive example: an LP is flagged as a honeypot if the token’s “buy tax” exceeds 50%, based on tax-related criteria from earlier research [25].
Profit validator. Profit-related activities reveal SLID characteristics (§IV-B3) from two insights: (i) the scam likelihood by evaluating realized profit, an essential element of all successful scams; (ii) the scam likelihood by evaluating the patterns of unrealized profits generated. Thus, we set both realized profit and unrealized profit as our metrics. Those values can be obtained through the collected DEX orders (more in §VI).
The LP owner’s realized profit can be computed by taking the difference between the total value of their investment in the LP (in USD) and the total value that has been withdrawn over the LP’s lifetime under the base token (detailed equation in Appendix A-C). We flag an LP $\mathcal{P}$ with its owner $\mathcal{O}$ having a positive total realized profit. Although not all projects that generate realized profits are scams, our observations in §IV-B3 show that all reported scams exhibit a positive realized profit.
The unrealized profit of the LP’s owner is another critical metric, especially when the SLID LP shows a significant growth in unrealized profit after the first month of deployment. The first-month unrealized profit represents the portion of profit (in the base tokens) belonging to the owner from the LP’s value after its initial month of deployment (detailed computing equation in Appendix A-C). In §IV-B3, we applied this unrealized profit metric for comparisons between SLID and rug pull scams, revealing a significant disparity between the average first-month unrealized profit ratios generated by the two groups (0.67% of the initial investment for rug pull and 156% of the initial investment for SLID LPs).
We set the threshold for unrealized profit in this heuristic to 0, due to the limited size of our initial SLID-like dataset from RQ1. The smallest computed unrealized profit in this dataset was $1.3e^{-14}$, but given the narrow sample, we acknowledge that this value may not reflect the broader range of real-world SLID scams. While restricting unrealized profit analysis to the first month could miss later-developing cases, applying a soft heuristic threshold that requires only a positive unrealized profit during this period allows us to capture highly relevant examples that align closely with known SLID characteristics.
Owner activity validator.  SLID scams prioritize a gradual approach by multiple small-proportion profit-taking actions over an extended period (§IV). We accordingly built Owner Activity Validator in the following steps:
  • $\bullet$
    excludes LPs where owners faithfully burn their initial (LP) tokens after deploying the LP (i.e., $\texttt{burned}(\texttt{LPT}_{\text{init}})=\text{true}$);
  • $\bullet$
    continue to detect the LPs where their owners retain control over the initial ($\texttt{LPT}_{\text{init}}$) tokens;
  • $\bullet$
    observe/analyse the owner’s activities in exploiting profits (whether $\mathcal{O}$ applies $\mathcal{W}$ or/and $\mathcal{S}$).
We measure their activities through two metrics: (i) Profit-taking order count, denoted by $c$, where $c=\texttt{count}(\mathcal{W})+\texttt{count}(\mathcal{S})$, accounts for the total number of DEX orders involved in profit extraction; and (ii) order impact, denoted by $I$, indicates the ratio of the $t^{th}$ profit-taking order size of value compared to the LP’s total value before the activity is executed. We define their thresholds as:
$c\geq{T_{\text{count}}},\,\text{where}\,I^{t}=\frac{q^{t}}{B^{t-1}}<T_{\text{impact}},\,\forall\ q\ \text{by}\ \mathcal{O},$ (1)
where $T_{\text{count}}$ and $T_{\text{impact}}$ are the thresholds (§IV) of the owner’s profit-taking orders counts and maximum impact on the LP.
We set $T_{\text{count}}=5$ as the lower threshold (the fewest profit-taking orders executed by $\mathcal{O}$) for flagging an LP as SLID (informed by our observations in \$IV). Meanwhile, we define $I^{t}$ as $I^{t}=\frac{q^{t}}{B^{t-1}}$, where $q^{t}$ represents the order size of any profit-taking order made by $\mathcal{O}$ at time $t$. $B^{t-1}$ is the LP size (total base token value in the LP) right before this order.
To reflect the characteristic of small-proportion orders, we set the impact threshold of all owners’ profit-taking orders to be less than 95% ($T_{\text{impact}}<0.95$) of LP size. Our analysis in §IV revealed that 95% was the lowest profit-taking impact recorded for a rug pull instance. It also ensures we capture a broader range of scams, including potential SLID scams that were unnoticed by rug pull detection heuristics. In comparison, for rug pull detection heuristics in [24, 39], only LPs with a 0.99 profit-taking impact are flagged.
Additionally, we decided not to incorporate adaptivity for all numerical thresholds because of the lack of ground truth. We focus on confirmed scams and mislabeled rug pulls. We acknowledge that attackers may adapt to the threshold and evade detection, but our model still lays a foundation.
[…]
Figure 5: Distribution of owner 1st month unrealized profit.
Figure 6: Distribution of owner 1st month profit-taking order count.
Figure 7: Distribution of the largest owner profit-taking order to the LP’s size.
Remark on heuristic validity. To evaluate the reliability of the heuristic and validate the selected thresholds and cut-offs, we conducted a series of empirical assessments:
  • $\bullet$
    We first conducted the false positive rate analyses to determine how often legitimate LPs are incorrectly flagged as malicious. We applied the heuristic to 100 randomly selected legitimate LPs, each associated with Ethereum tokens whose legitimacy is strongly supported by their creation or sponsorship by large, reputable funds or companies [58]. The heuristic correctly identified all non-SLID LPs as “not SLID”, without generating any false positives.
  • $\bullet$
    We then analyse the effectiveness of the selected numeric cut-offs and thresholds in distinguishing between legitimate and illegitimate LPs. To support this, we visualize their impacts in separating LP types (Figure 77). We use the aforementioned 100 legitimate LPs alongside illegitimate ones, including scams such as rug pulls, honeypots, and SLID, sourced from our initial DeFi scam lists (§IV-A). Beyond the initial honeypot and legitimate LP filtering, the remaining heuristics are designed to separate SLID from rug-pull LPs. These include thresholds based on: (i) unrealized profits after the first month; (ii) owner’s taking profit order count after the first month; and (iii) impact of taking order size on the LP.
    • $\circ$
      We demonstrate that the three thresholds are effective in distinguishing SLID from rug pulls. Each figure displays a red dashed line representing the cut-off value for the respective metric. In all cases, the box-and-whisker plots for SLID LPs lie entirely on one side of the threshold, while the rug-pull LPs fall entirely on the opposite side. When combined using a logical AND condition, these three filters achieve perfect separation between SLID and rug-pull LPs in our evaluation set. No rug-pull LP satisfies all three threshold conditions simultaneously, and all SLID LPs do, resulting in 100% accuracy for this subset.
[…]
VI Detecting SLID on Ethereum’s (RQ3)
@@
We conduct a large-scale scam detection by applying our heuristic (§V-B) to datasets (including LPs, user activities) collected from six leading Ethereum DEXs (§VI-A and §VI-B). We further propose an ML approach (§VI-C) to provide early warning of scams and detect borderline cases that the heuristic failed to capture (explanation in Appendix A-F1).
@@
Figure 8: Overall processes of our large-scale SLID detection on Ethereum.
VI-A Data Collection
@@
We collected longitudinal Ethereum’s LP data spanning over 67 months. We preprocessed the data to extract relevant features. We utilized both heuristic-based and ML algorithms to classify the LPs as SLID or non-SLID .
@@
Liquidity pool information.  We first collected all the existing LPs in Ethereum (denoted as PoolData) to be used for both detection methods. We utilized the well-known Geth client node [59] to retrieve all Ethereum block information, starting from block 6,627,000 (November 2, 2018, the date when the first LP was deployed) to block 20,081,867 (June 13, 2024, the time of writing this work). Then, we filtered only pools containing at least one well-established token (such as ETH, USDT, or USDC) as the pool’s legitimate base token. We avoid adding less-recognized and less-active pairs to the dataset.
@@
As a result, we gathered a total of 319,166 deployed LPs across the six largest DEXs on Ethereum [60] (Table I), including Uniswap [34], Curve Finance [12], Balancer [61], SushiSwap [13], PancakeSwap [62], and BancorSwap [63]. We extracted key features (Appendix A-H) for each harvested LP, including creation timestamps, token addresses involved in the pool, owner addresses, and the initial LP value in USD, using Chainalysis API [64].
DEX orders. Exchange activities and profit patterns within PoolData are essential for computing the profit-generating metrics (§V-B). We collected all types of exchange activities (via smart contract events) of six DEXs. Those activities include liquidity deposit, withdraw, and trade (buy/sell) between two tokens in each pool. Considering the data collection process operates independently of PoolData extraction, we specifically filtered the orders to include only those relevant to LPs in PoolData.
We collected 917M+ DEX activities across the deployed LPs (Table I). It accounts for 82.35% of the total Ethereum trading volume historically [65]. The resulting dataset, DEXData, contains rows where each entry represents an activity by either the LP owner or an investor. The dataset includes activities recorded up to July 13, 2024, where all pools in PoolData have at least one month of DEX activity history.
Smart contract security features.  Token security-related features (such as buy/sell tax, trading lock) from the paired token’s smart contracts are critical for applying the heuristic’s Honeypot Validator. We utilized the API provided by GoPlus [47] and De.Fi [42] to collect the token’s security characteristics indicative of honeypot behavior, including buy/sell tax, sell restriction, and more.
From the tokens (base and paired) in the collected LPs, we extracted their security features. These features were organized into a dataset, SecurityData. Each row represents a distinct token along with its complete set of security features.
VI-B Large-scale SLID Detection
We described how we apply the heuristic to detect Ethereum’s SLID scams. We began with data preprocessing and extracted the features for all LPs. We then conducted an in-depth analysis of the flagged SLID pools across multiple dimensions, including vitality, profit, and trend to create the first-ever SLID ground truth database.
Data preprocessing.  We preprocessed the three datasets (i.e., PoolData, DEXData, SecurityData) to enhance the feature set and prepare for the heuristic detection phase.
For each LP instance in PoolData, we extracted all related DEX activities from DEXData to compile its complete activity history. Profit-related metrics were then computed (equations in Appendix A-C) and added as new features to PoolData. These enriched metrics serve as inputs for the flagging process. Then, smart contract security features from SecurityData were mapped to corresponding LPs in PoolData.
Our preprocessing resulted in an updated dataset, called ExtendedPoolData, containing a comprehensive range of features (Table IV) required for heuristic detection.
Table I: Collected data on Ethereum.